OpenStudy (anonymous):

If g(t)=t^2-4, then the graph of g(t) crosses the x-axis when t equals? @whpalmer4

OpenStudy (whpalmer4):

Do you know the equation for the line that is the x-axis?

OpenStudy (anonymous):

what do you mean?

OpenStudy (whpalmer4):

the x-axis is a line, right?

OpenStudy (anonymous):

yea

OpenStudy (whpalmer4):

can you give me the equation of a line which is the same as the x-axis?

OpenStudy (whpalmer4):

it's just y = 0, right?

OpenStudy (whpalmer4):

so when \(g(t) = t^2-4\) crosses the \(x\) axis, what is the value of \(g(t)\)?

OpenStudy (anonymous):

um yea

OpenStudy (whpalmer4):

\[g(t) = 0\]when it crosses the \(x\) axis, right?

OpenStudy (anonymous):

yea

OpenStudy (whpalmer4):

So, we just need to solve \[g(t) = t^2-4 = 0\]for the value(s) of \(t\) that make that true.

OpenStudy (whpalmer4):

Any idea how to proceed with that?

OpenStudy (anonymous):

g(t)=t^2=4

OpenStudy (whpalmer4):

No, we need to solve \[g(t) = 0\]which means \[t^2-4=0\]because \[g(t) = t^2-4\]

OpenStudy (anonymous):

so is it -4?

OpenStudy (whpalmer4):

Is what -4?

OpenStudy (anonymous):

oh nvm

OpenStudy (whpalmer4):

The sum of -2 and -2? Yes, the sum of -2 and -2 is -4 :-)

OpenStudy (anonymous):

yea

OpenStudy (whpalmer4):

we're trying to find values of \(t\) that make \[t^2 -4 =0\]

OpenStudy (whpalmer4):

what if we were just trying to solve \[t -4 = 0\]what would you do?

OpenStudy (anonymous):

plus 4 on both sides

OpenStudy (anonymous):

which gets you t^2=4 and then square root both sides and you get t=2

OpenStudy (whpalmer4):

that's an answer. are there any others?

OpenStudy (anonymous):

-2

OpenStudy (whpalmer4):

very good. so let's check our work: \[g(t) = t^2- 4\]\[g(2) = (2)^2 -4 = 4-4=0\checkmark\]\[g(-2) = (-2)^2-4 = 4 - 4 = 0\checkmark\]

OpenStudy (anonymous):

wait i have a qusetion, why does it have to equal to 0 agian?

OpenStudy (whpalmer4):

we wanted the points where it crosses the x-axis. These are called "roots" or "zeros" or "solutions". The x-axis is equivalent to the line y = 0. If you want to find the points at which two functions intersect, you set the y values equal to each other and solve for the x values that make that true. \[y = 0\]\[y = t^2-4\]\[0=t^2-4\]\[t=\pm2\]

OpenStudy (whpalmer4):

Here's a graph.

OpenStudy (anonymous):

Ohh i see thank you, can i ask you some more questions tomorrow? my mom said i need to go to sleep

OpenStudy (whpalmer4):

I'll almost certainly be on tomorrow at some point. You know how to flag me down...I'll have a look later if I'm not on when you do so.

OpenStudy (anonymous):

when will you be on?

OpenStudy (whpalmer4):

Don't know.

OpenStudy (whpalmer4):

But I have a lot of work to do, so I'll probably be doing plenty of procrastinating by goofing around on OpenStudy :-)